how do i solve this problem,- The hundreds digit of a three-digit number is 2 more than the units digit. The digits of the original three-digit number are reversed, and the result is subtracted from the original three-digit number. What is the units digit of the result? Describe anything interesting you discover finding the solution. You need to try this out on at least three 3-digit numbers of your choosing.
Answer by josmiceli(9692)

Call the digits a,b, and c
The number is abc
given:
a = c + 2
abc - cba
The units digit will be c - (c + 2) ,
but to do this, you have to carry
10 over to the c, and you have
10 + c - (c + 2) = 8
check:
abc = 553
cba = 355
553 - 355 = 198
---------------
abc = 917
cba = 719
917 - 719 = 198
---------------
abc = 391
cba = 193
391 - 193 = 198
---------------
The result is always 198